# Accretion vs. Core-Filament Collision

### v1

@@ Abstract
Context. Recent millimeter and submillimeter observations have unveiled elongated and asymmetric structures around protostars. These structures, referred to as streamers, often exhibit coherent velocity gradients, seemingly indicating a directed gas flow towards the protostars. However, their origin and role in star formation remain uncertain.
Aims. A protostellar core Per-emb-2, located in Barnard 1, has a relatively large streamer with au, which is more prominent in emission from carbon chain molecules. We aim to unveil the formation mechanism of this streamer.
Methods. We conducted mapping observations towards Per-emb-2 using the Nobeyama 45-m telescope. We targeted carbon chain molecular lines such as CCS, HC3N, and HC5N at 45 GHz.
Results. Using astrodendro, we identified one protostellar and four starless cores, including three new detections, on the Herschel column density map. The starless and protostellar cores are more or less gravitationally bound. We discovered strong CCS and HC3N emissions extending from the north to the south, appearing to bridge the gap between the protostellar core and the starless core north of it. This bridge spans au with the velocities from 6.5 to 7.0 km s-1. The bridge has the velocity gradient opposite to the streamer. Thus, the streamer is unlikely to be connected to this bridge, suggesting that the streamer does not have an accretion origin.
Conclusions. We propose that a collision between a spherical core and the filament has shaped the density structure in this region, consequently triggering star formation within the head-tail-shaped core. In this core-filament collision (CFC) scenario, the collision appears to have fragmented the filament into two structures. The streamer is a bow structure, while the bridge is a remnant of the shock-compressed filament. Thus, we conclude that the Per-emb-2 streamer does not significantly contribute to the mass accumulation towards the protostar.
@@
stars: formation – ISM: clouds – ISM: structure – submillimeter: ISM1 Introduction
@@
Stars form from dense cores in molecular clouds. According to the standard scenario of star formation (Shu et al., 1987), a nearly spherical dense core undergoes gravitational contraction, forming a nearly axi-symmetric system consisting of a protostar and a rotationally-supported disk. However, recent observations have uncovered significant non-axisymmetric structures surrounding young protostellar systems, commonly referred to as streamers or spirals (Pineda et al., 2020; Sanhueza et al., 2019; Chen et al., 2021; Valdivia-Mena et al., 2022; Hsieh et al., 2023; Olguin et al., 2023). Hereafter, we will use the term ”streamer” to refer to such structures for the sake of simplicity. These structures often exhibit coherent velocity patterns, interpreted as infall towards the central protostar (Pineda et al., 2020; Sanhueza et al., 2021). Understanding the role of streamers in mass loading towards central stars is crucial for unveiling the complexities of stellar mass accumulation process.
@@
Numerical studies have proposed various mechanisms for streamer formation. For example, in many numerical simulations of cluster formation, the intermediate-density structure connected with dense cores and sink particles, which resemble the streamers, are often observed and these structures partly play a role in mass accretion onto the central protostellar system (e.g., Bonnell & Bate, 2006; Nakamura & Li, 2007; Kuffmeier et al., 2017; Padoan et al., 2020). On the other hand, Tu et al. (2024) demonstrated that streamer-like structures could arise from local turbulent compression within magnetized parent cores, without the need for external accretion of ambient gas. In this case, the protostellar system does not have to have structures connected with the inter-core medium, and streamers can be generated through the accretion process within the natal core. Therefore, the accretion of ambient gas is likely to be a minor process for supplying the gas, and the stellar mass is roughly determined by the parent core mass. Yano et al. (2024) proposed that streamers could form through dense core collisions (DCC, see also Kinoshita & Nakamura, 2022). They estimated the frequencies of DCC in nearby star-forming molecular clouds and found that the typical dense core experiences at least a collision in its lifetime, particularly in clustered environments. In this DCC model, the compressed layer created between two cores evolves into streamers and spirals, creating the rotating envelope around the protostellar system. In this case, the streamer contributes to supply additional mass into the central protostar, but the mass accretion occurs temporarily, depending on the conditions such as the impact parameter and collision speed. However, these mechanisms await further elucidation based on observational evidence.
@@
To shed light on the streamer formation, we conducted observations towards Per-emb-2, a core with a streamer towards the central protostar (Pineda et al., 2020). It is located in the southern part of Barnard 1 in the Perseus molecular cloud at a distance of 300 pc from the Sun (Zucker et al., 2019). Gas within the streamer appears to exhibit a velocity gradient towards the central protostar [from 7.5 km s-1 (far side) to 7.0 km s-1 (protostar)], and the streamer is particularly prominent in emissions from chemically early-phase molecules like CCS and HC3N as observed with the NOEMA interferometer. Since these molecules steeply decrease with time at the high-density environment, Pineda et al. (2020) claim that the gas of the streamer is supplied from outside via accretion. Very recently, Taniguchi et al. (2024) also claimed that the ambient gas may be provided from the northern part of the protostar.
@@ 2 Observations and data
We carried out the on-the-fly (OTF) mapping observations towards the protostellar core Per-emb-2 using the Z45 receiver of the NRO 45m telescope (Nakamura et al., 2014) in the 2022–2023 season. Some details of the observations are presented in the Appendix. The mapped area is determined such that the protostar is located at the centre of the box (Fig. 1a).
The target lines are CCS (), HC3N (), and HC5N (). The parameters are summarised in Table 2. The telescope has a full-width at half-maximum (FWHM) beam size of 38″ at 43 GHz. The standard chopper wheel method was used to convert the output signal into the antenna temperatures () and corrected for the atmospheric attenuation. The main beam efficiency of Z45 is at 45 GHz. The main-beam temperature was calculated as . We adopted a spheroidal function as a gridding convolution function to calculate the intensity at each grid point of the final cube data with a spatial grid size of 15″ and a frequency resolution of 3.81 kHz which corresponds to km s-1 at 45 GHz. The final effective angular resolution of the map was 49″, corresponding to 15000 au, or 0.07 pc at the distance of 300 pc (Zucker et al., 2019).
We downloaded the high resolution H2 column density and dust temperature maps obtained by the Herschel Gould Belt Survey (André et al., 2010; Pezzuto et al., 2021). The angular resolution of the maps are 18″.
@@
1 | 3:32:19.13 | 30:51:44.67 | 0.508 | 0.0148 | 1.28 | 5.4 | 12.7 | 0.37 | 1.422 (0.565) | starless/449 |
2 | 3:32:16.36 | 30:51:08.40 | 0.398 | 0.0077 | 1.25 | 29.6 | 12.6 | 0.41 | 1.014 (0.377) | starless/new |
3 | 3:32:18.50 | 30:49:50.63 | 9.723 | 0.0884 | 1.97 | 0.5 | 12.1 | 0.84 | 1.077 (0.176) | protostellar/447 |
4 | 3:32:26.94 | 30:48:21.45 | 0.200 | 0.0130 | 0.64 | 3.1 | 13.8 | – | – (1.265) | starless/new |
@@ 3.1 Dense core identification and their dynamical states
In Fig. 1a, we present the Herschel column density map of the observed area. We applied astrodendro (Rosolowsky et al., 2008) to this image and identified five cores, four cores fully covered by our observation box (see the Appendix B.1 for more details). The boundaries of the identified cores are outlined in red. Their physical quantities are summarised in Table 1. In the same area, Pezzuto et al. (2021) identified three cores using getsources: a protostellar core (HGBS 447) and two starless cores (HGBSs 442, 449). Our core Nos. 1 and 3 correspond to HGBS 442 and 447, respectively. The cores Nos. 2, 4 and 5 are starless and newly detected. We could not identify HGBS 449, which has the lowest column density. The central protostellar core (No. 3) has a mass of . According to Pineda et al. (2020), the protostellar core has a systemic velocity of km s-1 and has a streamer with a length of au (Pineda et al., 2020).
@@
From the CCS cube data, we derived the core’s CCS velocity dispersion () and then converted it to the intrinsic velocity dispersion () of gas with a mean molecular weight of to assess their dynamical state. See the Appendix B.2 for the details of the analysis. The cores Nos. 1, 2, 3, and 5 have , where the virial parameter was computed under the assumption of centrally-condensed () sphere. The protostellar core has a significantly-small Bonnor-Ebert ratio with 0.18, but is close to unity, where the Bonnor-Ebert ratio is the ratio between the Bonnor-Ebert critical mass of the core and the observed mass. The core No. 4 does not have significant CCS emission, and we could not measure the velocity dispersion, but the Bonnor-Ebert ratio of the core No. 4 ( 1.3) suggests it is close to bound. We also calculate the external pressures exerted on the core surfaces, and found that they play a minor role except for No. 5, where we calculated the core’s external pressure () from the ambient density and ambient gas intrinsic velocity dispersion measured in the branch just below the corresponding core. The branch is an intermediate structure identified by astrodendro. The core No.5 is dynamically compressed by the ambient pressure and may be hard to create from spontaneous gravitational fragmentation for which cores become self-gravitating. The other cores are close to be in dynamical equilibrium.
@@
For the protostellar core, the column density distribution is not symmetric, but exhibiting a head-tail shape. The centre of gravity determined from this distribution exhibits a significant offset from the position of the protostar itself (see Fig. 8). This suggests that the streamer does not point directly towards the gravity centre of the protostellar core. If its curved structure is generated by the angular momentum of the infalling gas, a large ( au) rotating envelope should form around the protostar and connect to the streamers. However, there is currently no observational evidence for such a large structure (e.g., Tobin et al., 2016; Pineda et al., 2020). Therefore, it is unlikely that the streamer is driven by gravitational infall from outside the core.
@@
In Fig. 1, we present the velocity integrated maps of CCS, HC3N, and HC5N alongside with the Herschel column density map. In the CCS integrated intensity map, we recognise the two prominent CCS elongated features: an elongated feature, stretching from north to south, bridging the gap between the starless cores Nos. 1&2 and the protostellar core No. 3, while another compact emission is positioned to the west of the protostellar core.
@@
The distribution of HC3N emission appears to resemble that of CCS, with strong concentrations at the southern periphery of the starless cores Nos. 1&2. Another significant peak is evident immediately above the protostar, with a fainter peak located to the west, coinciding with a local maximum in CCS intensity. HC5N, exhibiting the weakest intensity among the three lines, has a peak at a similar position of the second peak of HC3N, or the northern portion of the protostellar core.
@@
Figs. 2 & 7 display the velocity channel maps of CCS, HC5N, and HC3N. The starless cores Nos. 1&2 appear associated with relatively strong compact emission around a velocity of 6.5 km s-1.
@@
The CCS emission within the velocity range of 6.5–6.7 km s-1 exhibits a fragmented filamentary structure, extending along a line from the northeast to the southwest which coincides with the large-scale filament seen in the H2 column density image (see Fig. 8). An elongated structure is discernible in the southeast corner of the CCS map, though the peak is faint in the H2 map (see Fig. 1).
@@
At a velocity of 6.6–7.0 km s-1, the elongated structure observed in CCS gradually connects with the central protostellar core, forming a bridge. A similar bridging structure is also observed in the corresponding velocity channel for HC5N. Overall, these emissions demonstrate a weak velocity gradient spanning from 6.5 km s-1 to 7.0 km s-1 towards the protostar.
3.3 Mismatch in velocity structure between bridge and streamer
@@
Based on the observations detailed in the previous section, it becomes evident that the protostellar core No. 3 and the northern starless cores Nos.1&2 exhibit a connection characterized by a bridge structure. In Fig. 3, we present the CCS position-velocity plot tracing the 0.2 pc bridge structure.
@@
The protostellar core No. 3 exhibits a velocity of 7 km s-1, indicating a velocity disparity between the two cores, bridged by an intermediate velocity region. The velocity gradient along the bridge is estimated to be 0.5 km s-1/0.2 pc 2.5 km s-1 pc-1. The existence of the intermediate-velocity gas bridging the two structures with different velocities suggests a dynamical interaction between the cores, a similar feature to the larger-scale cloud-cloud collisions (e.g., Nakamura et al., 2012; Nguyen-Lu’o’ng et al., 2013; Fukui et al., 2021; Kinoshita et al., 2021). The enhanced CCS emission observed could potentially stem from the colliding region, where CCS formation may be accelerated due to increased density. Both CCS and HC3N exhibit critical densities around cm-3, implying that while the shocked layer may possess larger density, the column density might not be sufficiently high to be recognised in the 2D Herschel image.
@@
The streamer seemingly follows the cm-2 contour of the protostellar core and is spatially overlapped with the distribution of the CCS emission (see Fig. 2), but the velocities are very different. The streamer has much steeper velocity gradient in the opposite direction to that of the bridge (see the right panel of Fig. 3). Therefore, we conclude that the bridge is not a structure connected to the streamer.

### v2

# Accretion versus core-filament collision

@@ Abstract
Context. Recent millimetre and sub-millimetre observations have unveiled elongated and asymmetric structures around protostars. These structures, referred to as streamers, often exhibit coherent velocity gradients, seemingly indicating a directed gas flow towards the protostars. However, their origin and role in star formation remain uncertain.
Aims. The protostellar core Per-emb-2, located in Barnard 1, has a relatively large streamer of au that is more prominent in emission from carbon-chain molecules. We aim to unveil the formation mechanism of this streamer.
Methods. We conducted mapping observations towards Per-emb-2 using the Nobeyama 45 m telescope. We targeted carbon-chain molecular lines such as CCS, HC3N, and HC5N at 45 GHz.
Results. Using astrodendro, we identified one protostellar and four starless cores, including three new detections, on the Herschel column density map. The starless and protostellar cores are more or less gravitationally bound. We discovered strong CCS and HC3N emissions extending from the north to the south, appearing to bridge the gap between the protostellar core and the starless core to its north. This bridge spans au with velocities of 6.5 to 7.0 km s-1. The velocity gradient of the bridge is opposite that of the streamer. Thus, the streamer is unlikely to be connected to the bridge, suggesting that the former does not have an accretion origin.
Conclusions. We propose that a collision between a spherical core and the filament has shaped the density structure in this region, consequently triggering star formation within the head-tail-shaped core. In this core-filament collision scenario, the collision appears to have fragmented the filament into two structures. The streamer is a bow structure, while the bridge is a remnant of the shock-compressed filament. Thus, we conclude that the Per-emb-2 streamer does not significantly contribute to the mass accumulation towards the protostar.
@@
stars: formation - ISM: clouds - ISM: structure - submillimeter: ISM1 Introduction
@@
Stars form from dense cores in molecular clouds. According to the standard scenario of star formation (Shu et al., 1987), a nearly spherical dense core undergoes gravitational contraction, forming a nearly axi-symmetric system consisting of a protostar and a rotationally supported disk. However, recent observations have uncovered significant non-axisymmetric structures surrounding young protostellar systems, commonly referred to as streamers or spirals (Pineda et al., 2020; Sanhueza et al., 2019; Chen et al., 2021; Valdivia-Mena et al., 2022; Hsieh et al., 2023; Olguin et al., 2023). Hereafter, we use the term ‘streamer’ to refer to such structures for the sake of simplicity. These structures often exhibit coherent velocity patterns, interpreted as infall towards the central protostar (Pineda et al., 2020; Sanhueza et al., 2021). Understanding the role of streamers in mass loading towards central stars is crucial for unveiling the complexities of the stellar mass accumulation process.
@@
Numerical studies have proposed various mechanisms for streamer formation. For example, in many numerical simulations of cluster formation, intermediate-density structures connected to dense cores and sink particles, which resemble the streamers, are often observed, and these structures play a role in mass accretion onto the central protostellar system (e.g. Bonnell & Bate, 2006; Nakamura & Li, 2007; Kuffmeier et al., 2017; Padoan et al., 2020). On the other hand, Tu et al. (2024) demonstrate that streamer-like structures could arise from local turbulent compression within magnetised parent cores, without the need for external accretion of ambient gas. In this case, the protostellar system does not need to have structures connected to the inter-core medium, and streamers can be generated through the accretion process within the natal core. Therefore, the accretion of ambient gas likely plays only a minor role in supplying the gas, and the stellar mass is roughly determined by the parent core mass. Yano et al. (2024) propose that streamers could form through dense core collisions (DCCs; see also Kinoshita & Nakamura, 2022). They estimated the frequencies of DCCs in nearby star-forming molecular clouds and found that the typical dense core experiences at least one collision in its lifetime, particularly in clustered environments. In this DCC model, the compressed layer created between two cores evolves into streamers and spirals, creating a rotating envelope around the protostellar system. In this case, the streamer supplies additional mass to the central protostar, but the mass accretion occurs only temporarily and depends on conditions such as the impact parameter and collision speed. However, observational evidence is needed to further elucidate these mechanisms.
@@
To shed light on streamer formation, we conducted observations towards Per-emb-2, a core with a streamer towards the central protostar (Pineda et al., 2020). It is located in the southern part of Barnard 1 in the Perseus molecular cloud at a distance of 300 pc from the Sun (Zucker et al., 2019). Gas within the streamer appears to exhibit a velocity gradient towards the central protostar – from 7.5 km s-1 (far side) to 7.0 km s-1 (protostar) – and the streamer is particularly prominent in emissions from chemically early-phase molecules such as CCS and HC3N, as observed with the Northern Extended Millimetre Array (NOEMA) interferometer. Since the abundance of these molecules decreases steeply with time in high-density environments, Pineda et al. (2020) claim that the gas of the streamer is supplied from outside via accretion. Very recently, Taniguchi et al. (2024) suggested that the ambient gas may be provided from the northern part of the protostar.
@@ 2 Observations and data
We carried out the on-the-fly mapping observations towards the protostellar core Per-emb-2 using the Z45 receiver of the Nobeyama Radio Observatory (NRO) 45m telescope (Nakamura et al., 2014) in the 2022–2023 season. Details of the observations are presented in Appendix A. The mapped area was determined such that the protostar is located at the centre of the box (Fig. 1a).
The target lines are CCS (), HC3N (), and HC5N (). The parameters are summarised in Table 2. See Nakamura et al. (2024) for the accurate rest frequency of CCS (). The telescope has a full width at half maximum beam size of 38″ at 43 GHz. The standard chopper wheel method was used to convert the output signal to the antenna temperatures () and to correct for the atmospheric attenuation. The main beam efficiency of Z45 is at 45 GHz. The main-beam temperature was calculated as . We adopted a spheroidal function as a gridding convolution function to calculate the intensity at each grid point of the final cube data with a spatial grid size of 15″ and a frequency resolution of 3.81 kHz, which corresponds to km s-1 at 45 GHz. The final effective angular resolution of the map was 49″, corresponding to 15000 au, or 0.07 pc at the distance of 300 pc (Zucker et al., 2019).
We downloaded the high-resolution H2 column density and dust temperature maps obtained by the Herschel Gould Belt Survey (André et al., 2010; Pezzuto et al., 2021). The angular resolution of the maps is 18″.
@@
| 1 | 3:32:19.13 | 30:51:44.67 | 0.508 | 0.0148 | 1.28 | 5.4 | 12.7 | 0.37 | 1.422 (0.565) | starless/449 |
| 2 | 3:32:16.36 | 30:51:08.40 | 0.398 | 0.0077 | 1.25 | 29.6 | 12.6 | 0.41 | 1.014 (0.377) | starless/new |
| 3 | 3:32:18.50 | 30:49:50.63 | 9.723 | 0.0884 | 1.97 | 0.5 | 12.1 | 0.84 | 1.077 (0.176) | protostellar/447 |
| 4 | 3:32:26.94 | 30:48:21.45 | 0.200 | 0.0130 | 0.64 | 3.1 | 13.8 | – | – (1.265) | starless/new |
@@ 3.1 Dense core identification and their dynamical states
In Fig. 1a we present the Herschel column density map of the observed area. We applied astrodendro (Rosolowsky et al., 2008) to this image and identified five cores, four of which are fully within our observation box (see Appendix B.1 for more details). The boundaries of the identified cores are outlined in red. Their physical quantities are summarised in Table 1. In the same area, Pezzuto et al. (2021) identified three cores using getsources: a protostellar core (HGBS 447) and two starless cores (HGBSs 442 and 449). Our cores 1 and 3 correspond to HGBS 442 and 447, respectively. Cores 2, 4, and 5 are starless and newly detected. We could not identify HGBS 449, which has the lowest column density. The central protostellar core (core 3) has a mass of . According to Pineda et al. (2020), the protostellar core has a systemic velocity of km s-1 and has a streamer with a length of au (Pineda et al., 2020).
@@
From the CCS cube data, we derived the core’s CCS velocity dispersion () and then converted it to the intrinsic velocity dispersion () of gas with a mean molecular weight of to assess its dynamical state (see Appendix B.2 for the details of the analysis). Cores 1, 2, 3, and 5 have ; the virial parameter was computed under the assumption of a centrally condensed () sphere. The protostellar core has a significantly small Bonnor-Ebert ratio of 0.18, but its is close to unity (the Bonnor-Ebert ratio is the ratio between the Bonnor-Ebert critical mass of the core and the observed mass). Core 4 does not have significant CCS emission, and we could not measure the velocity dispersion, but its Bonnor-Ebert ratio ( 1.3) suggests it is close to bound. We also calculated the external pressures exerted on the core surfaces. We find that they play a minor role except for core 5, for which we calculated the core’s external pressure () from the ambient density () and the ambient gas intrinsic velocity dispersion () measured in the branch just below the corresponding core. The branch is an intermediate structure identified by astrodendro. Core 5 is dynamically compressed by the ambient pressure and likely did not form via spontaneous gravitational fragmentation, a process that leads to cores becoming self-gravitating. The other cores are close to being in dynamical equilibrium.
@@
For the protostellar core, the column density distribution is not symmetric: it exhibits a head-tail shape. The centre of gravity determined from this distribution is significantly offset from the position of the protostar itself (see Fig. 8). This suggests that the streamer does not point directly towards the gravity centre of the protostellar core. If its curved structure is generated by the angular momentum of the infalling gas, a large rotating envelope ( au) should form around the protostar and connect to the streamers. However, there is currently no observational evidence of such a large structure (e.g. Tobin et al., 2016; Pineda et al., 2020). Therefore, it is unlikely that the streamer is driven by gravitational infall from outside the core.
@@
In Fig. 1 we present the velocity-integrated maps of CCS, HC3N, and HC5N alongside the Herschel column density map. In the CCS integrated intensity map, we recognise the two prominent CCS elongated features: (1) an elongated feature, stretching from north to south and bridging the gap between the starless cores 1 and 2 and the protostellar core 3, and (2) a compact emission positioned to the west of the protostellar core.
@@
The distribution of HC3N emission appears to resemble that of CCS, with strong concentrations at the southern periphery of the starless cores 1 and 2. Another significant peak is evident immediately above the protostar, and a fainter peak is located to the west, coinciding with a local maximum in CCS intensity. HC5N, which exhibits the weakest intensity of the three lines, has a peak at a similar position as the second peak of HC3N, or the northern portion of the protostellar core.
@@
Figures 2 and 7 display the velocity channel maps of CCS, HC5N, and HC3N. The starless cores 1 and 2 appear to be associated with relatively strong compact emission around a velocity of 6.5 km s-1.
@@
The CCS emission within the velocity range 6.5–6.7 km s-1 exhibits a fragmented filamentary structure, extending along a line from the north-east to the south-west that coincides with the large-scale filament seen in the H2 column density image (see Fig. 8). An elongated structure is discernible in the south-east corner of the CCS map, though the peak is faint in the H2 map (see Fig. 1).
@@
At a velocity of 6.6–7.0 km s-1, the elongated structure observed in CCS gradually connects with the central protostellar core, forming a bridge. A similar bridging structure is also observed in the corresponding velocity channel for HC5N. Overall, these emissions demonstrate a weak velocity gradient that spans from 6.5 km s-1 to 7.0 km s-1 towards the protostar.
3.3 Mismatch in velocity structure between the bridge and streamer
@@
From the observations detailed in the previous section, it becomes evident that the protostellar core 3 and the northern starless cores 1 and 2 exhibit a connection characterised by a bridge structure. In Fig. 3 we present a CCS position-velocity plot that traces the 0.2 pc bridge structure.
@@
The protostellar core 3 exhibits a velocity of 7 km s-1, indicating a velocity disparity between the two cores bridged by an intermediate-velocity region. The velocity gradient along the bridge is estimated to be 0.5 km s-1/0.2 pc 2.5 km s-1 pc-1. The existence of the intermediate-velocity gas bridging the two structures with different velocities suggests a dynamical interaction between the cores, a feature similar to larger-scale cloud-cloud collisions (e.g. Nakamura et al., 2012; Nguyen-Lu’o’ng et al., 2013; Fukui et al., 2021; Kinoshita et al., 2021). The enhanced CCS emission observed could potentially stem from the colliding region, where CCS formation may be accelerated due to increased density. Both CCS and HC3N exhibit critical densities of around cm-3, implying that while the shocked layer may have a higher density, the column density might not be sufficiently high to be recognised in the 2D Herschel image.
@@
The streamer seemingly follows the cm-2 contour of the protostellar core and spatially overlaps with the distribution of the CCS emission (see Fig. 2), but the velocities are very different. The streamer has a much steeper velocity gradient in the opposite direction to that of the bridge (see the right panel of Fig. 3). Therefore, we conclude that the bridge is not a structure connected to the streamer.